# ToM-LM: Delegating Theory of Mind Reasoning to External Symbolic Executors in Large Language Models

### v2

@@ 1 Introduction
Theory of Mind (ToM) refers to an individual’s ability to conceptualize and attribute mental states to others, therefore comprehending and representing how others perceive, interpret, and represent the surrounding world [16, 2, 17]. This ability is essential for individuals to understand and reason about others’ beliefs, intentions, emotions, obligations, and so on [16, 5, 17]. For Large Language Models (LLMs), in most scenarios, they are required to interact with users or other agents and to understand their needs to give effective and appropriate responses. Therefore, discovering and enhancing their ToM reasoning ability is necessary and critical. Furthermore, it is worth noting that reasoning about knowledge has been a concern in symbolic philosophy and logic for many decades [4] and certainly in computer science to model security and game theory [1, 6, 7]. To a large extent, it is also the case that most of the machine learning models today have not really supported epistemic modeling and learning. Therefore, the idea that capabilities classically dominated in symbolic computing could be used in the context of LLMs is interesting and presents intriguing possibilities.
Previous studies have demonstrated the emergence of ToM reasoning ability in LLMs [10, 13, 8, 11]. Nevertheless, most of these studies adopt the ToM evaluation problems originally designed for assessing human ToM reasoning ability, which could be widely exposed to the internet, leading to possible data leakage to the training processes of LLMs, and because of the simplicity of these ToM problems, in order to study whether LLMs can handle more complex ToM reasoning ability, new benchmarks are needed [12, 20]. Recent work has developed a new benchmark called MindGames [20] aiming to give a more complex ToM benchmark specifically designed for evaluating the ToM reasoning ability of LLMs. This benchmark also indicates that despite increases in model size, LLMs still struggle with complex ToM reasoning tasks [20].
@@
Within this approach, an LLM is first given a ToM problem, and then it is asked to convert the problem represented in the natural language to the symbolic formulation. Subsequently, the generated symbolic formulation is then executed by the SMCDEL model checker to perform the ToM reasoning in a transparent and verifiable way and give out the final answer. However, our investigations reveal that this strategy alone does not adequately enhance the ToM reasoning ability of LLMs. Therefore, we employ fine-tuning [^1] We fine-tune the chosen LLM, gpt-3.5-turbo, through OpenAI Fine-tuning Platform with pairs of natural language and symbolic formulation. More details are discussed later in this paper. to further optimize this approach, aiming to substantially improve the ToM reasoning ability of LLMs. The overview of ToM-LM is shown in Figure 1.
Our study demonstrates that ToM-LM improves the ToM reasoning ability of LLMs significantly in terms of accuracy and Area under the ROC Curve (AUC), compared to all the constructed baselines. The contributions of this study are outlined as follows:
@@ 2 Related Work
In addition to the discussion in the previous section, we reiterate the following related works.
@@ 2.1 ToM In LLMs
Previous works have discussed more around whether ToM ability has emerged in LLMs [10, 8, 18, 21, 3]. Some works also try in various ways to improve the ToM reasoning ability of LLMs. As in [13], the various in-context learning methods are utilized such as Let’s think step by step [9] and the chain-of-thought [23] to enhance the ToM reasoning ability of LLMs. Although this approach brings up an easy and flexible way to improve LLMs’ ToM reasoning ability, we still hold the view that dynamic and complex ToM problems can not be simply represented as a chain-of-thought reasoning process and a more appropriate and robust way to handle them is needed. Thus, we propose an approach to use symbolic formulation to represent a general way to capture the ToM reasoning process and use external symbolic executors to perform the reasoning.
@@ 3.1.1 Model Fine-tuning
During this stage, we fine-tune the LLM, specifically the gpt-3.5-turbo, using our newly constructed Symbolic Formulation Generation Prompting Fine-tuning Dataset [^2] We explain Symbolic Formulation Generation Prompting Fine-tuning Dataset later in Section 3.2. which consists of pairs of ToM problems represented in both natural language and symbolic formulation. The fine-tuning was being processed on the OpenAI Fine-tuning Platform with 3 epochs. This process aims to improve the performance of the generation of symbolic formulations.
@@ 3.1.2 Problem Prompting
Subsequently, a ToM problem and a one-shot in-context example are constructed as a prompt and given to the fine-tuned LLM to instruct it to give an appropriate symbolic formulation. Figure 2 shows the template used for prompting in which the texts within the braces serve as placeholders to be replaced with actual example or problem contents. The placeholders beginning with "example" indicate they are for the one-shot in-context example while the placeholders starting with "problem" mean they are for the ToM problem.
Figure 2: Symbolic Formulation Generation Prompting Template
@@ 3.1.3 Problem Formalization
The fine-tuned LLM is then expected to generate the symbolic formulation to appropriately represent the problem described in the natural language. An example symbolic formulation is illustrated in Figure 3. The generated symbolic formulations should conform to this structure and are anticipated to be executable.
@@ 3.2 Dataset
We select MindGames [20] as our evaluation and fine-tuning dataset. It is a dataset designed specifically for evaluating the ToM reasoning ability of the LLMs. Each item in the dataset includes a premise which describes the ToM problem situation and context, a hypothesis which describes a situation needing LLMs to judge its validity, a Boolean label which is the ground truth of the validity of the hypothesis, and a corresponding symbolic formulation to the premise and hypothesis.
For example, as illustrated in Figure 3, the premise describes a scenario involving four agents, each drawing a face unrevealed card, and someone may reveal their card to others. The hypothesis to be validated based on this given scenario is "Cara can now know whether Conrad picked a red card". The corresponding symbolic formulation representation of the premise and hypothesis is also given. To briefly explain [^3] Refer to https://github.com/jrclogic/SMCDEL for more syntax details., "VARS 1,2,3,4" defines four atomic propositions which represent the status of each agent, such as that "2" represents "Conrad picked a red card". "LAW Top" indicates that all possible Boolean states of these statuses should be considered. "OBS" simply denotes which agent observes which status given in "VARS", such as that "Agentc:1" means Cara observes the status that "Vasiliki picked a red card". The "VALID?" line queries an agent’s belief about a certain status, in which "!" means a public announcement to all agents and "|" is a logical OR, such as that "!(1|2|3|4)" means "It is publicly announced that someone picked a red card". It represents and encapsulates the main part of the premise and hypothesis.
Each problem in MindGames can be categorized as either a true-belief or false-belief problem. LLMs are expected to handle them correctly to show their ToM reasoning ability.
@@
Additionally, we randomly sample 150 items from the train set of MindGames dataset to form as the fine-tuning dataset. While maintaining the same data, the dataset is further preprocessed in two separate ways to suit different fine-tuning purposes: (1) Direct Prompting Fine-tuning Dataset, wherein each problem is given to the Direct Prompting Template as shown in Figure 4 to generate the appropriate prompt, and the corresponding target ground truth Boolean label is stringified as "TRUE" or "FALSE" and it is prepared specifically for the Direct Prompting With Fine-tuning baseline; (2) Symbolic Formulation Generation Prompting Fine-tuning Dataset, in which each problem is given to Symbolic Formulation Generation Prompting Template as shown in Figure 2 to generate the appropriate prompt, and the corresponding target is not the Boolean label but the ground truth symbolic formulation, which is prepared for ToM-LM.
@@ 3.3 Baselines
It stands for Direct Prompting Without Fine-tuning. In this setting, we directly ask the base model to answer the problem in "TRUE", "FALSE", or "I DON’T KNOW", with a one-shot in-context example provided. Answering in "TRUE" means it considers the hypothesis is valid in the given situation, "FALSE" means it thinks the hypothesis is invalid, and "I DON’T KNOW" means it has no idea about the validity of the hypothesis. The prompting template is illustrated in Figure 4.
Figure 4: Direct Prompting Template
@@ 3.3.2 SFGP
It stands for Symbolic Formulation Generation Prompting Without Fine-tuning. In this setting, the workflow is exactly as same as ToM-LM but the model is not fine-tuned.
@@ 3.3.3 DPFT
It stands for Direct Prompting With Fine-tuning. In this setting, the workflow is exactly as same as the DP baseline but the model is fine-tuned based on the Direct Prompting Fine-tuning Dataset.
@@ 4.1.1 Execution Rate
The execution rate of baselines in Direct Prompting setting are almost around 100%, which means the LLM can generate available answers at most of time in this setting. Instead, in the case of Symbolic Formulation Generation Prompting setting, the execution rate is 78% without fine-tuning but can be high up to 94.5% in ToM-LM.
@@ 4.1.2 Accuracy
As demonstrated in Table 1, ToM-LM reaches up to 91% accuracy which significantly surpasses all the baselines. It achieves the 33% improvement over the DP baseline, the 42% increases compared to the SFGP baseline, the 15% enhancement over the DPFT baseline.
@@ 5.1 Results Analysis
The results demonstrate that ToM-LM has successfully enhanced the ToM reasoning ability of LLMs, surpassing the other three baselines in terms of both accuracy, which reaches as high as 91%, and AUC which attains to 0.94.
From the results of the DP baseline, the accuracy is 58% and AUC is 0.58, meaning although the LLM itself may possess a modest degree of ToM reasoning ability, it still struggles with it and its answers are highly dependent on random guessing. From the SFGP baseline, although AUC is shown with a marginal improvement by 0.02 and reaches to 0.6, indicating that it modestly enhances the performance of the LLM against random guessing, the gain comes at the cost of reduced accuracy and a highly increased incidence of non-executable symbolic formulations.
As anticipated, the DPFT baseline markedly surpasses both of the DP and SFGP baseline in terms of both accuracy and AUC. When comparing it to our approach, ToM-LM shows significant improvement over it with both of accuracy and AUC. In addition, within the same fine-tuning settings, we also observe that the training loss of the DPFT baseline has a violent oscillation and does not converge at the end but the training loss of the ToM-LM has a very slighter oscillation and it converges at the end of fine-tuning.
@@
5.2 Implications
Our study introduces a novel approach to externalize the ToM reasoning process to make it transparent and explainable and to significantly improve the ToM reasoning ability of LLMs in cases of true-belief and false-belief tasks. We hope it can potentially provide a new perspective on the feasibility of externalizing and delegating other aspects of ToM reasoning to external executors.
Furthermore, given that the essence of ToM involves understanding others in multi-agent contexts, this study is also anticipated to provide valuable insights for real-world applications, including multi-agent collaboration, autonomous driving, and human-machine interactions.
@@ 5.3 Limitations and Future Work
ToM could and is supposed to encompass a broad range of aspects, such as emotions, goals, intentions, desires, and beyond, besides the foundation of handling false-belief and true-belief. Due to the expressive limitations of the current external executor, our study focuses solely on a specific aspect of ToM, handling false-belief and true-belief. Nevertheless, the framework established here has the potential to be generalized to address various ToM-related tasks. Future research should explore additional components of ToM using external executors and may also aim to develop a more expressive and capable external executor or language to handle a diverse array of ToM tasks.
Moreover, our research primarily utilizes gpt-3.5-turbo. Future studies should investigate whether this approach is replicable with language models in smaller parameter sizes or explore it in more powerful models to determine if the approach can consistently and efficiently enhance ToM reasoning ability.

### v3

@@ 1 Introduction
Theory of Mind (ToM) refers to an individual’s ability to attribute mental states to others [16, 2, 17]. This ability is essential for individuals to understand and reason about others’ beliefs, intentions, emotions, obligations, and so on [16, 5, 17]. For Large Language Models (LLMs), in most scenarios, they are required to interact with users or other agents and to understand their needs to give effective and appropriate responses. Therefore, discovering and enhancing their ToM reasoning ability is necessary and critical. Furthermore, it is worth noting that reasoning about knowledge has been a concern in symbolic philosophy and logic for many decades [4] and certainly in computer science to model security and game theory [1, 6, 7]. To a large extent, it is also the case that most of the machine learning models today have not really supported epistemic modeling and learning. Therefore, the idea that capabilities classically dominated in symbolic computing could be used in the context of LLMs is interesting and presents intriguing possibilities.
Previous studies have demonstrated the emergence of ToM reasoning ability in LLMs [10, 13, 8, 11]. Nevertheless, most of these studies adopt the ToM evaluation problems originally designed for assessing human ToM reasoning ability, which could be widely exposed to the internet, leading to possible data leakage to the training processes of LLMs [12, 20]. In addition, because of the simplicity of these ToM problems, in order to study whether LLMs can handle more complex ToM reasoning ability, new benchmarks are needed [12, 20]. Recent work has developed a new benchmark called MindGames [20] aiming to give a more complex ToM benchmark specifically designed for evaluating the ToM reasoning ability of LLMs. This benchmark also indicates that despite increases in model size, LLMs still struggle with complex ToM reasoning tasks [20].
@@
In our approach, an LLM is first given a ToM problem, and then it is asked to convert the problem represented in the natural language to the symbolic formulation. Subsequently, the generated symbolic formulation is then executed by the SMCDEL model checker to perform the ToM reasoning in a transparent and verifiable way and give out the final answer. However, our investigations reveal that this strategy alone does not adequately enhance the ToM reasoning ability of LLMs. Therefore, we employ fine-tuning [^1] We fine-tune the chosen LLM, gpt-3.5-turbo, through OpenAI Fine-tuning Platform with pairs of natural language and symbolic formulation. More details are discussed later in this paper. to further optimize this approach, aiming to substantially improve the ToM reasoning ability of LLMs. The overview of ToM-LM is shown in Figure 1.
Our study demonstrates that ToM-LM significantly improves the ToM reasoning ability of LLMs in terms of accuracy and Area under the ROC Curve (AUC), compared to all constructed baselines. The contributions of this study are outlined as follows:
@@ 2.1 ToM In LLMs
Previous works have discussed the emergence of ToM abilities in LLMs [10, 8, 18, 21, 3]. Some works also try in various ways to improve the ToM reasoning ability of LLMs. As in [13], the various in-context learning methods are utilized such as Let’s think step by step [9] and the chain-of-thought [23] to enhance the ToM reasoning ability of LLMs. Although this approach brings up an easy and flexible way to improve LLMs’ ToM reasoning ability, we still hold the view that dynamic and complex ToM problems can not be simply represented as a chain-of-thought reasoning process and a more appropriate and robust way to handle them is needed. Thus, we propose an approach to use symbolic formulation to represent a general way to capture the ToM reasoning process and use external symbolic executors to perform the reasoning.
@@ 3.1.1 Model Fine-tuning
During this stage, we fine-tune the LLM, specifically the gpt-3.5-turbo, using our newly constructed Symbolic Formulation Generation Prompting Fine-tuning Dataset [^2] Symbolic Formulation Generation Prompting Fine-tuning Dataset is explained in Section 3.2. which consists of pairs of ToM problems represented in both natural language and symbolic formulation. The fine-tuning was being processed on the OpenAI Fine-tuning Platform with 3 epochs. This process aims to improve the performance of the generation of symbolic formulations.
@@ 3.1.2 Problem Prompting
Subsequently, a ToM problem and a one-shot in-context example are constructed as a prompt and given to the fine-tuned LLM to instruct it to give an appropriate symbolic formulation [^3] Details of the prompting template are described in Appendix 0.A.2..
@@ 3.1.3 Problem Formalization
The fine-tuned LLM is then expected to generate the symbolic formulation to appropriately represent the problem described in the natural language. The generated symbolic formulations are anticipated to be executable [^4] An example of the symbolic formulation is given in Appendix 0.B..
@@ 3.2 Dataset
We select MindGames [20] as our evaluation and fine-tuning dataset. It is a dataset designed specifically for evaluating the ToM reasoning ability of the LLMs. Each item [^5] An example is given in Appendix 0.B. in the dataset represents a ToM problem including a premise which describes the ToM problem situation and context, a hypothesis which describes a situation needing LLMs to judge its validity, a Boolean label which is the ground truth of the validity of the hypothesis, and a corresponding symbolic formulation to the premise and hypothesis. Each problem in MindGames can be categorized as either a true-belief or false-belief problem. LLMs are expected to handle them correctly to show their ToM reasoning ability.
@@
Additionally, we randomly sample 150 items from the MindGames train set for fine-tuning. The dataset is preprocessed in two ways for different purposes: (1) Direct Prompting Fine-tuning Dataset, in which each problem is formatted using the Direct Prompting Template [^6] Details of the prompting template are described in Appendix 0.A.1. to generate prompts prepared for the DPFT baseline; (2) Symbolic Formulation Generation Prompting Fine-tuning Dataset, in which each problem is formatted using the Symbolic Formulation Generation Prompting Template [^7] Details of the prompting template are described in Appendix 0.A.2. to generate prompts prepared for ToM-LM.
@@ 3.3 Baselines
It stands for Direct Prompting Without Fine-tuning. In this setting, we directly ask the base model to answer the problem in "TRUE", "FALSE", or "I DON’T KNOW", with a one-shot in-context example provided. Answering in "TRUE" means it considers the hypothesis is valid in the given situation, "FALSE" means it thinks the hypothesis is invalid, and "I DON’T KNOW" means it has no idea about the validity of the hypothesis.
@@ 3.3.2 SFGP
It stands for Symbolic Formulation Generation Prompting Without Fine-tuning. In this setting, the workflow is the same as ToM-LM but the model is not fine-tuned.
@@ 3.3.3 DPFT
It stands for Direct Prompting With Fine-tuning. In this setting, the workflow is the same as the DP baseline but the model is fine-tuned based on the Direct Prompting Fine-tuning Dataset.
@@ 4.1.1 Execution Rate
The execution rates of baselines in Direct Prompting setting are almost around 100%, which means the LLM can generate available answers at most of time in this setting. Instead, in the case of Symbolic Formulation Generation Prompting setting, the execution rate is 78% without fine-tuning but can be high up to 94.5% in ToM-LM.
@@ 4.1.2 Accuracy
As demonstrated in Table 1, ToM-LM reaches up to 91% accuracy which significantly surpasses all the baselines. It achieves the 33% improvement over the DP baseline, the 42% increase compared to the SFGP baseline, the 15% enhancement over the DPFT baseline.
@@ 5.1 Results Analysis
The results demonstrate that ToM-LM has successfully enhanced the ToM reasoning ability of LLMs, surpassing the other three baselines in terms of both accuracy, which reaches as high as 91%, and AUC which attains 0.94.
From the results of the DP baseline, the accuracy is 58% and AUC is 0.58, meaning although the LLM itself may possess a modest degree of ToM reasoning ability, it still struggles with it and its answers are highly dependent on random guessing. From the SFGP baseline, although AUC is shown with a marginal improvement by 0.02 and reaches 0.6, indicating that it modestly enhances the performance of the LLM against random guessing, the gain comes at the cost of reduced accuracy and a highly increased incidence of non-executable symbolic formulations.
As anticipated, the DPFT baseline markedly surpasses both of the DP and SFGP baseline in terms of both accuracy and AUC. When comparing it to our approach, ToM-LM shows significant improvement over it with both of accuracy and AUC. In addition, within the same fine-tuning settings, we also observe that the training loss of the DPFT baseline has a violent oscillation and does not converge at the end but the training loss of the ToM-LM has a very slighter oscillation and it converges at the end of fine-tuning [^8] Loss curves for both DPFT and ToM-LM are given in Appendix 0.C..
@@ 5.2 Limitations and Future Work
While ToM is supposed to encompass a broad range of aspects, such as emotions, goals, intentions, desires, and beyond, our study only focuses on handling false-belief and true-belief of ToM due to the expressive limitations of the current external executor. Nevertheless, this approach has the potential to be generalized to handle other ToM tasks. Future research should explore additional components of ToM using external executors and may also aim to develop a more expressive and capable external executor or language to handle a diverse array of ToM tasks.
@@ 6 Conclusion
Appendix 0.A Prompting Templates
0.A.1 Direct Prompting Template
Figure 3 shows the template used for prompting to directly generate the prediction of the answer. The texts within the braces serve as placeholders to be replaced with actual example or problem contents. The placeholders beginning with "example" indicate that they are for the one-shot in-context ToM example, while the placeholders "context" and "hypothesis" are for the ToM problem.
Figure 3: Direct Prompting Template
0.A.2 Symbolic Formulation Generation Prompting Template
Figure 4 demonstrates the template used for prompting to generate the prediction of the symbolic formulation. The texts within the braces serve as placeholders to be replaced with actual example or problem contents. The placeholders beginning with "example" indicate that they are for the one-shot in-context ToM example while the placeholders starting with "problem" mean they are for the ToM problem.
Figure 4: Symbolic Formulation Generation Prompting Template
Appendix 0.B Problem Example
An example of a ToM problem consisting of a premise, a hypothesis, and a symbolic formulation given in MindGames [20] is illustrated in Figure 5.
The premise describes a scenario involving four agents, each drawing a face unrevealed card, and someone may reveal their card to others. The hypothesis to be validated based on this given scenario is "Cara can now know whether Conrad picked a red card". The corresponding symbolic formulation representation of the premise and hypothesis is also given. To briefly explain [^9] Refer to https://github.com/jrclogic/SMCDEL for more syntax details., "VARS 1,2,3,4" defines four atomic propositions which represent the status of each agent, such as that "2" represents "Conrad picked a red card". "LAW Top" indicates that all possible Boolean states of these statuses should be considered. "OBS" simply denotes which agent observes which status given in "VARS", such as that "Agentc:1" means Cara observes the status that "Vasiliki picked a red card". The "VALID?" line queries an agent’s belief about a certain status, in which "!" means a public announcement to all agents and "|" is a logical OR, such as that "!(1|2|3|4)" means "It is publicly announced that someone picked a red card". It represents and encapsulates the main part of the premise and hypothesis.
Appendix 0.C Fine-tuning Loss Curves
0.C.1 Loss Curve of Direct Prompting With Fine-tuning
As the loss curve illustrated in Figure 6, the training loss of the DPFT baseline is high with the range of loss between 0 and 9 as shown in the y-axis, while also exhibits severe oscillations and failing to converge by the end of training.
Figure 6: Loss Curve of Direct Prompting With Fine-tuning
0.C.2 Loss Curve of ToM-LM
As shown in Figure 7, the training loss of the ToM-LM is low with the range of loss between 0 and 0.9 and converges at the end of training.
Figure 7: Loss Curve of ToM-LM